OpenStudy (anonymous):

URGENT! HELP PLEASE, I'LL REWARD YOU solve f(x) =x^2-4x+3 show work

OpenStudy (anonymous):

@mathmate

OpenStudy (anonymous):

@mathlover2014

OpenStudy (lifeisadangerousgame):

Solve for x?

OpenStudy (anonymous):

for f(-2)

OpenStudy (anonymous):

@LifeIsADangerousGame

OpenStudy (lifeisadangerousgame):

Oh, okay. So when it says solve for f(x) = x^2-4x+3 when f(-2) it means plug in -2 for x in the equation. So: f(-2) = -2^2 - 4(-2) + 3 =

OpenStudy (anonymous):

im so confused. i was taught to substitute f(x) with y

OpenStudy (anonymous):

can you help me please and show the work? cause i know after that i have to switch out the variables

OpenStudy (anonymous):

and i have to turn this in within the next 5 minutes

OpenStudy (bossvideogamer21):

i will help too

OpenStudy (anonymous):

thank you!

OpenStudy (lifeisadangerousgame):

Okay: f(x) = x^2-4x+3 f(-2) = (-2)^2-4(-2)+3 f(-2) = 4 + 8 + 3 f(-2) = 15

OpenStudy (bossvideogamer21):

what does the sign right next to the x mean

OpenStudy (lifeisadangerousgame):

Where at?

OpenStudy (bossvideogamer21):

in the question it says x^

OpenStudy (bossvideogamer21):

so what does this ^ mean?

OpenStudy (lifeisadangerousgame):

f(x) IS y, but in this case you would not put -2 for y because this is not that type of problem. ^ means exponent. So x^2 is the same as \[x ^{2}\]

OpenStudy (anonymous):



OpenStudy (bossvideogamer21):

oh thank u

OpenStudy (anonymous):

NP

OpenStudy (lifeisadangerousgame):

Pro tip: Often times, when you plug in -2^2 or something similar into a calculator, the calculator will read it is -(2^2) and will give you the wrong answer. To prevent this, plug in (-2)^2, this forces the calculator to realize you are squaring a negative number.

OpenStudy (bossvideogamer21):

here is the real answer:

OpenStudy (bossvideogamer21):

ax^2+bx+c x^2-4x+3 (x-3)(x-1)=0 x=3, x=1 when y=0 The graph of the parabola has a minimum because a is positive. (3+1)/2=2 so y will be at a minimum when x=2 f(x)=y=2^2-(4*2)+3 =4-8+3

OpenStudy (bossvideogamer21):

may i have my reward now?

OpenStudy (bossvideogamer21):

after all i did answer ur question

OpenStudy (mathmate):

@rosamartinez "solve" and "evaluate/calculate" have different meanings in mathematics. For your information, to "solve" means to find the unknown given and equation, for example, solve 5x+3=18, then we "solve" by isolating x on the left, and find that x=3. When we are given all information to calculate the value of a function, we say "evaluate" f(-2), since x=-2, and the function is given/defined. For the given case, we can "solve" f(x)=x^2-4x+3=0 (if =0 was given) in which case f(x)=(x-3)(x-1)=0, or x=3 or x=1. or "evaluate" f(-3), then f(-3)=(-3)^2-4(-3)+3=9+12+3=24.